# Beyond Pairwise Learning-To-Rank At Airbnb

### v1

@@ Abstract.
There are three fundamental asks from a ranking algorithm: it should scale to handle a large number of items, sort items accurately by their utility, and impose a total order on the items for logical consistency. But here’s the catch—no algorithm can achieve all three at the same time. We call this limitation the SAT theorem for ranking algorithms. Given the dilemma, how can we design a practical system that meets user needs? Our current work at Airbnb provides an answer, with a working solution deployed at scale.
@@ 1. Introduction
Search ranking orders a corpus of items to maximize some combination of utilities for the searcher. It’s the tool we rely on daily to cut through endless choices—whether it’s finding some information, a meal, a job, a vacation spot, or a place to stay. At the heart of search ranking is the learning-to-rank (LTR) algorithm, which learns from past searcher decisions to improve future results. A straightforward way to learn from past examples is to frame it as a binary classification task, where items that satisfied the searchers’ needs are treated as positive, the remaining choices that were presented to the searcher are labeled negative, and a model is trained to distinguish between the two. However a lot of industrial scale applications use, not the simple binary classification, but a variant of the formulation known as pairwise LTR.
The seminal paper that introduced the Bradley-Terry-Luce model and analyzed pairwise LTR appeared in 1957  (Ford, 1957), with neural network applications to the problem emerging in a 1994 paper  (Bromley et al., 1993). Since then, pairwise LTR has found adoption at major internet companies  (Sculley et al., 2009), (Senthilkumar et al., 2024), (Cao et al., 2007), (Sharma and Yan, 2013), including Airbnb  (Haldar et al., 2019). What has kept this decades old area of research active and thriving? It’s the increasing availability of data, advances in computational power, and the growing business value of improved rankings. This paper presents one step forward in this long narrative arc, highlighting recent advancements in ranking at Airbnb that go beyond traditional pairwise LTR.
We start by examining the properties that make pairwise LTR scalable and effective. The formalism introduced in Section 3 is a stepping stone for the sections ahead, where we address the limitations of pairwise LTR that impact its accuracy. We make improvements in two steps, first considering interactions between two listings at a time in Section 4, then in Section 5 scaling up to interactions across all listings. Section 6 extends the framework from optimizing for a single objective, to optimizing for multiple objectives. We cite related work embedded in context of the discussion throughout the paper. The final section covers experiments performed at Airbnb to compare the frameworks developed, and their outcomes. But first, the SAT theorem.
@@ 3. Pairwise LTR: A Brief Overview
$l_{booked}$ : features of the booked listing
$l_{notBooked}$ : features of the listing not booked
@@
(1) $loss=-log(\theta(f(l_{booked})-f(l_{notBooked})))$
@@
The place where pairwise LTR misses out is accuracy. Here’s why. Ordering the listings by $f(\cdot)$ is equivalent to ordering them by the pairwise booking probabilities. The Bradley-Terry equation links these pairwise booking probabilities to the individual (pointwise) booking probabilities $P(l_{a})$ and $P(l_{b})$:
@@
For a more precise understanding of pairwise LTR’s limitations, let’s focus on the properties of the pairwise logit described in Eq 2. If we take three listings, $l_{a}$, $l_{b}$, and $l_{c}$, it’s easy to show that:
(5) $logit(a,b)+logit(b,c)=logit(a,c)$
@@
This restriction on the logit space prevents pairwise LTR from accurately modeling booking probabilities for users. Let’s break it down with an example. Imagine two listings, $l_{x}$ and $l_{y}$, that are identical apartments in the same building, but $l_{x}$ is \$10 cheaper. Naturally, most searchers would prefer $l_{x}$, so $logit(l_{x},l_{y})$ would have a high positive value, i.e., $logit(l_{x},l_{y})\gg 0$.
Now, add a third listing, $l_{z}$, that’s quite different in terms of location, price, and amenities. The values of $logit(l_{x},l_{z})$ and $logit(l_{y},l_{z})$ would probably be similar because users who favor $l_{x}$ over $l_{z}$ would also favor $l_{y}$ over $l_{z}$. This means $logit(l_{x},l_{z})\approx logit(l_{y},l_{z})$. If we apply collinearity, we’d get
$logit(l_{x},l_{y})+logit(l_{y},l_{z})\approx logit(l_{y},l_{z})$
which implies $logit(l_{x},l_{y})\approx 0$. This contradicts the earlier result of $logit(l_{x},l_{y})\gg 0$, proving that collinearity doesn’t apply in this case. Figure 1 illustrates why pairwise comparisons can’t always fit into a 1D logit space.
Figure 1. The relation $logit(l_{x},l_{y})\gg 0\land logit(l_{x},l_{z})\approx logit(l_{y},l_{z})$ cannot be captured in a 1D space. It is represented by the triangle inequality $logit(l_{x},l_{y})+logit(l_{y},l_{z})>logit(l_{x},l_{z})$.
@@ 4. True-Pairwise LTR
(6) $loss=-log(\theta(g(l_{bk},l_{nbk})))-log(1-\theta(g(l_{nbk},l_{bk})))$
which is the symmetric binary cross entropy loss for predicting that $l_{bk}$ is preferred over $l_{nbk}$.
@@ 4.1. Total Order
To solve this problem given $N$ listings, for each listing $l_{i}$, we collect the $(N-1)$ scores $g(l_{i},l_{1}),g(l_{i},l_{2}),\dots,g(l_{i},l_{n})$. Given that we have established anti-commutativity for $g(\cdot)$, we can apply a simpler version of the generalized Bradley-Terry model (Eq 10 from  (Huang et al., 2006)):
(9) $score(l_{i})=\frac{1}{1+\sum_{j\neq i,j=1}^{N}e^{-g(l_{i},l_{j})}}$
These scores assign each listing a real value and are invariant to how the listings are ordered. Since real-valued scores naturally fulfill the requirements for total order, they can now be used with standard sorting algorithms like Quicksort or Mergesort. Figure 2(b) illustrates how a listing is scored in the true-pairwise LTR framework.
(a) True-pairwise training, optimizing the loss from Eq 6.
(b) True-pairwise scoring utilizing Eq 9.
Figure 2. True-pairwise LTR
Our approach can be contrasted with the true-pairwise LTR method proposed in  (Nardini et al., [n. d.]), where the $(N-1)$ scores are averaged for a listing. By leveraging the generalized Bradley-Terry model, we achieve a higher accuracy of booking probability due to the performance bounds provided in  (Huang et al., 2006). The two approaches are compared under experimental results.
@@ 4.2. Scalability
With $N$ listings, true-pairwise LTR requires $N*(N-1)/2$ unique evaluations of $g(\cdot)$, one for each pair of listings. In contrast, pairwise LTR only needs $N$ evaluations. Because of its quadratic complexity, we use true-pairwise LTR only in the final stage of a multi-stage ranking system, after the number of listings has been reduced to $O(10)$. This leads to a hybrid system where the first-stage ranker is still owned by pairwise LTR. Figure 3 outlines the complexity levels across different ranking stages. For more on algorithms suited to the final ranking stage, see  (Liu et al., 2022).
@@ 5. All-Pairwise LTR
Bivariate true-pairwise LTR improves the accuracy of univariate pairwise LTR by accounting for interactions between two listings being compared. Building on this idea, we introduce a multivariate LTR framework that considers how each listing interacts with all other listings at once. This approach better aligns with user behavior described in choice theory (Tversky, 1972) and leads to improved accuracy of booking probability predictions.
@@
  • Calculating the $(N-1)$ pairwise logits using the true-pairwise LTR model $g(\cdot)$.
  • Combining these pairwise logits into one final score using the generalized Bradley-Terry model (Eq 9).
In all-pairwise LTR, we replace both these steps by a single model. This model not only takes the listing being scored as input, but the remaining $(N-1)$ listings as well to compute interactions.
The generalized Bradley-Terry model, used in the second step of true-pairwise LTR, lacks adaptability since it doesn’t have learnable parameters. To address this, we replace it with an MLP, named the all-pairwise logit network (APLN). During training, APLN generates logits, which we pair for booked and non-booked listings, and then optimize the loss in Eq 1.
Next we turn our attention to the features that should feed the APLN network. To efficiently capture how a listing interacts with the other $(N-1)$ listings, we create two types of features:
@@
If superiority-similarity features are directly fed into the APLN network, input permutation invariance cannot be guaranteed. To see this, consider a single neuron of the input layer of APLN. Let the $i$th input (e.g. superiority over $l_{i}$) be $x_{i}$, the $i$th weight be $w_{i}$, and $b$ be the layer bias. The neuron’s output, $x_{i}*w_{i}+b$, changes if $l_{i}$ is reordered. To make the neuron’s output invariant to input order, we make the $i$th weight $w_{i}$ depend on $l_{i}$. This dependency is modeled through a function $\phi(l_{i})$, implemented as an MLP. The modified neuron output, written as $x_{i}*\phi(l_{i})+b$, is now permutation invariant.
An alternative way to achieve permutation invariance, proposed in  (Ai et al., 2019), involves generating and aggregating permutations of item subsets. In our approach, we avoid the computational complexity of creating permutations. The superiority-similarity features, along with the layer imposing input permutation invariance, make up the all pairwise feature network (APFN). Construction of superiority features is outlined in Figure 4(a), while Figure 4(b) illustrates the similarity features. The overall all-pairwise LTR architecture is shown in Figure 5.
(a) Feature computing the superiority of the given listing $l_{i}$ over the remainder of listings $l_{1},l_{2},\dots l_{N}$. $f(\cdot)$ is the pairwise LTR first-stage ranker represented by Eq 2. $\theta(\cdot)$ is the sigmoid function. $\phi(\cdot)$ is an MLP which takes the listing features as input. $\beta$ is the bias term.
@@
Figure 4. Superiority and similarity features of APFN
Figure 5. The all-pairwise architecture. Superiority features refers to Figure 4(a). Similarity features refers to Figure 4(b). Together they form the all pairwise feature network (APFN). The APFN feeds into the all pairwise logit network (APLN), an MLP. To produce the final logit, APLN acts as a residual to the logit produced by the pairwise LTR first-stage ranker $f(\cdot)$.
We designed APFN to learn how the scored listing interacts with the other $N-1$ listings. An alternative to APFN is to use an attention network to construct the features, as explored in  (Pasumarthi et al., 2020),  (Wei et al., 2020) and  (Pang et al., 2020). This gives us an off-the-shelf solution which is easier to get started with. In the experimental results section to follow, we compare how these methods stack up against each other.
@@
Another interesting point of comparison for all-pairwise LTR is with neural networks that implement differentiable sorting functions, such as  (Grover et al., 2019) and  (Prillo and Eisenschlos, 2020). These too are $O(N^{2})$ complexity solutions capable of ranking listings, but their focus is to replace traditional sorting algorithms like Quicksort with a neural network. In contrast, the goal of all-pairwise LTR is to improve ranking by learning interactions between listings, while still using Quicksort or Mergesort to perform the final sorting.
A frequent stumbling block is the apparent resemblance between all-pairwise LTR and its distant cousin, listwise LTR (Cao et al., 2007), a univariate model that optimizes the softmax cross-entropy loss. Since listwise LTR is univariate, it cannot account for interactions between listings. Similarly, methods that directly optimize NDCG  (Qin et al., 2010),  (Jagerman et al., 2022) also rely on univariate models, leaving them unable to model listing-to-listing interactions.
@@ 6. Optimizing Multiple Objectives
A popular solution to the problem is multi-task learning ((Ren et al., 2023), (Tan et al., 2023)), where a model predicts multiple objectives through different heads and combines them using weights. This approach makes the model more complex, since it has to handle multiple labels and balance dense and sparse ones. Particularly in our case, true-pairwise and all-pairwise LTR fit only one of the objectives, the booking probability, as searchers naturally compare various listings in search results. These methods aren’t suitable for predicting trip ratings, requiring a separate model to handle that task.
@@
We discard the multi-task and multi-model approaches to solve the multiple objectives puzzle. Instead, we unify the objectives by adjusting the training loss. For each training example consisting of a pair of listings $l_{booked}$ and $l_{notBooked}$, we compute a weight:
(10) $\omega_{total}=1+\omega_{tripQuality}(l_{booked})$
where a booking get a base weight of $1$, and trip quality adds to it depending on the reviews of $l_{booked}$. The loss is calculated as:
(11) $loss=-log(\theta(f(l_{booked})-f(l_{notBooked})))*\omega_{total}$

### v2

@@ Abstract.
Abstract: There are three fundamental asks from a ranking algorithm: it should scale to handle a large number of items, sort items accurately by their utility, and impose a total order on the items for logical consistency. But here’s the catch—no algorithm can achieve all three at the same time. We call this limitation the SAT theorem for ranking algorithms. Given the dilemma, how can we design a practical system that meets user needs? Our current work at Airbnb provides an answer, with a working solution deployed at scale.
@@ 1. Introduction
Search ranking orders a corpus of items to maximize some combination of utilities for the searcher. It’s the tool we rely on daily to cut through endless choices—whether it’s finding some information, a meal, a job, a vacation spot, or a place to stay. At the heart of search ranking is the learning-to-rank (LTR) algorithm, which learns from past searcher decisions to improve future results. A straightforward way to learn from past examples is to frame it as a binary classification task, where items that satisfied the searchers’ needs are treated as positive, the remaining choices that were presented to the searcher are labeled negative, and a model is trained to distinguish between the two. However, many industrial scale applications use not the simple binary classification, but a variant of the formulation known as pairwise LTR.
The seminal paper that introduced the Bradley-Terry-Luce model and analyzed pairwise LTR appeared in 1957  (Ford, 1957), with neural network applications to the problem emerging in a 1994 paper  (Bromley et al., 1993). Since then, pairwise LTR has found adoption at major internet companies  (Sculley et al., 2009) (Senthilkumar et al., 2024) (Cao et al., 2007) (Sharma and Yan, 2013), including Airbnb  (Haldar et al., 2019). What has kept this decades-old area of research active and thriving? It’s the increasing availability of data, advances in computational power, and the growing business value of improved rankings. This paper presents one step forward in this long narrative arc, highlighting recent advancements in ranking at Airbnb that go beyond traditional pairwise LTR.
We start by examining the properties that make pairwise LTR scalable and effective. The formalism introduced in Section 3 is a stepping stone for the sections ahead, where we address the limitations of pairwise LTR that impact its accuracy. We make improvements in two steps, first considering interactions between two listings at a time in Section 4, then in Section 5 scaling up to interactions across all listings. Section 6 extends the framework from optimizing for a single objective to optimizing for multiple objectives. We cite related work embedded in context of the discussion throughout the paper. The final section covers experiments performed at Airbnb to compare the frameworks developed, and their outcomes. But first, the SAT theorem.
@@ 3. Pairwise LTR: A Brief Overview
$l_{\text{booked}}$ : features of the booked listing
$l_{\text{notBooked}}$ : features of the listing not booked
@@
(1) $\text{loss}=-\log(\theta(f(l_{\text{booked}})-f(l_{\text{notBooked}})))$
@@
The place where pairwise LTR misses out is accuracy. Here’s why. Ordering the listings by $f(\cdot)$ is equivalent to ordering them by the pairwise booking probabilities. The Bradley-Terry equation  (wiki:bradleyterry) links these pairwise booking probabilities to the individual (pointwise) booking probabilities $P(l_{a})$ and $P(l_{b})$:
@@
For a more precise understanding of pairwise LTR’s limitations, let’s focus on the properties of the pairwise logit described in Equation 2. If we take three listings, $l_{a}$, $l_{b}$, and $l_{c}$, it’s easy to show that:
(5) $\text{logit}(a,b)+\text{logit}(b,c)=\text{logit}(a,c)$
@@
This restriction on the logit space prevents pairwise LTR from accurately modeling booking probabilities for users. Let’s break it down with an example. Imagine two listings, $l_{x}$ and $l_{y}$, that are identical apartments in the same building, but $l_{x}$ is \$10 cheaper. Naturally, most searchers would prefer $l_{x}$, so $\text{logit}(l_{x},l_{y})$ would have a high positive value, i.e., $\text{logit}(l_{x},l_{y})\gg 0$.
Now add a third listing, $l_{z}$, that’s quite different in terms of location, price, and amenities. The values of $\text{logit}(l_{x},l_{z})$ and $\text{logit}(l_{y},l_{z})$ would be similar because users who favor $l_{x}$ over $l_{z}$ would also favor $l_{y}$ over $l_{z}$. This means $\text{logit}(l_{x},l_{z})\approx\text{logit}(l_{y},l_{z})$. If we apply collinearity, we’d get
$\text{logit}(l_{x},l_{y})+\text{logit}(l_{y},l_{z})\approx\text{logit}(l_{y},l_{z})$
which implies $\text{logit}(l_{x},l_{y})\approx 0$. This contradicts the earlier result of $\text{logit}(l_{x},l_{y})\gg 0$, proving that collinearity doesn’t apply in this case. Figure 1 illustrates why pairwise comparisons can’t always fit into a 1D logit space.
Figure 1. The relation $\text{logit}(l_{x},l_{y})\gg 0\land\text{logit}(l_{x},l_{z})\approx\text{logit}(l_{y},l_{z})$ cannot be captured in a 1D space. It is represented by the triangle inequality $\text{logit}(l_{x},l_{y})+\text{logit}(l_{y},l_{z})>\text{logit}(l_{x},l_{z})$.
@@ 4. True-Pairwise LTR
(6) $\text{loss}=-\log(\theta(g(l_{bk},l_{nbk})))-\log(1-\theta(g(l_{nbk},l_{bk})))$
which is the symmetric binary cross-entropy loss for predicting that $l_{bk}$ is preferred over $l_{nbk}$.
@@ 4.1. Total Order
To solve this problem given $N$ listings, for each listing $l_{i}$, we collect the $N-1$ scores $g(l_{i},l_{1}),g(l_{i},l_{2}),\dots,g(l_{i},l_{n})$. Given that we have established anti-commutativity for $g(\cdot)$, we can apply a simpler version of the generalized Bradley-Terry model (Equation 10 from  (Huang et al., 2006)):
(9) $\text{score}(l_{i})=\frac{1}{1+\sum_{j\neq i,j=1}^{N}e^{-g(l_{i},l_{j})}}$
These scores assign each listing a real value and are invariant to how the listings are ordered initially. Since real-valued scores naturally fulfill the requirements for total order, they can now be used with standard sorting algorithms like Quicksort or Mergesort. Figure 2(b) illustrates how a listing is scored in the true-pairwise LTR framework.
(a) True-pairwise training, optimizing the loss from Equation 6.
(b) True-pairwise scoring utilizing Equation 9.
Figure 2. True-pairwise LTR.
Our approach can be contrasted with the true-pairwise LTR method proposed in  (Nardini et al., [n. d.]), where the $N-1$ scores are averaged for a listing. By leveraging the generalized Bradley-Terry model, we achieve a higher accuracy of booking probability due to the performance bounds provided in  (Huang et al., 2006). The two approaches are compared under experimental results.
@@ 4.2. Scalability
With $N$ listings, true-pairwise LTR requires $N(N-1)/2$ unique evaluations of $g(\cdot)$, one for each pair of listings. In contrast, pairwise LTR only needs $N$ evaluations of $f(\cdot)$. Because of its quadratic complexity, we use true-pairwise LTR only in the final stage of a multi-stage ranking system, after the number of listings has been reduced to $O(10)$. This leads to a hybrid system where the first-stage ranker is still owned by pairwise LTR. Figure 3 outlines the complexity levels across different ranking stages. For more on algorithms suited to the final ranking stage, see  (Liu et al., 2022).
@@ 5. All-Pairwise LTR
Bivariate true-pairwise LTR improves the accuracy of univariate pairwise LTR by accounting for interactions between the two listings being compared. Building on this idea, we introduce a multivariate LTR framework that considers how each listing interacts with all the other listings at once. This approach better aligns with the user behavior described in choice theory (Tversky, 1972) and leads to improved accuracy of booking probability predictions.
@@
  1. (1)
    Calculating the $N-1$ pairwise logits using the true-pairwise LTR model $g(\cdot)$.
  2. (2)
    Combining these pairwise logits into one final score using the generalized Bradley-Terry model (Equation 9).
In all-pairwise LTR, we replace both these steps by a single model. This model not only takes the listing being scored as input, but the remaining $N-1$ listings as well to compute interactions.
The generalized Bradley-Terry model, used in the second step of true-pairwise LTR, lacks adaptability since it doesn’t have learnable parameters. To address this, we replace it with an MLP named the all-pairwise logit network (APLN). During training the APLN generates logits, which we pair for booked and non-booked listings, and then optimize the loss in Equation 1.
Next we turn our attention to the features that should feed the APLN. To efficiently capture how a listing interacts with the other $N-1$ listings, we create two types of features:
@@
If superiority-similarity features are directly fed into the APLN, input permutation invariance cannot be guaranteed. To see this, consider a single neuron of the input layer of the APLN. Let the $i$th input (e.g. superiority over $l_{i}$) be $x_{i}$, the $i$th weight be $w_{i}$, and $b$ be the layer bias. The neuron’s output, $x_{i}*w_{i}+b$, changes if $l_{i}$ is reordered. To make the neuron’s output invariant to input order, we make the $i$th weight $w_{i}$ depend on $l_{i}$. This dependency is modeled through a function $\phi(l_{i})$, implemented as an MLP. The modified neuron output, written as $x_{i}*\phi(l_{i})+b$, is now permutation invariant.
An alternative way to achieve permutation invariance, proposed in  (Ai et al., 2019), involves generating and aggregating permutations of item subsets. In our approach, we avoid the computational complexity of creating permutations. The superiority-similarity features, along with the layer imposing input permutation invariance, make up the all-pairwise feature network (APFN). Construction of superiority features is outlined in Figure 4(a), while Figure 4(b) illustrates the similarity features. The overall all-pairwise LTR architecture is shown in Figure 5.
(a) Feature computing the superiority of the given listing $l_{i}$ over the remainder of listings $l_{1},l_{2},\dots l_{N}$. $f(\cdot)$ is the pairwise LTR first-stage ranker represented by Equation 2. $\theta(\cdot)$ is the sigmoid function. $\phi(\cdot)$ is an MLP which takes the listing features as input. $\beta$ is the bias term.
@@
Figure 4. Superiority and similarity features of APFN.
Figure 5. The all-pairwise architecture. Superiority features refers to Figure 4(a). Similarity features refers to Figure 4(b). Together they form the all-pairwise feature network (APFN). The APFN feeds into the all-pairwise logit network (APLN), an MLP. To produce the final logit, APLN acts as a residual to the logit produced by the pairwise LTR first-stage ranker $f(\cdot)$.
We designed APFN to learn how the scored listing interacts with the other $N-1$ listings. An alternative to APFN is to use an attention network to construct the features, as explored in  (Pasumarthi et al., 2020),  (Wei et al., 2020), and  (Pang et al., 2020). This gives us an off-the-shelf solution which is easier to get started with. In the experimental results section to follow, we compare how these methods stack up against each other.
@@
Another interesting point of comparison for all-pairwise LTR is with neural networks that implement differentiable sorting functions, such as  (Grover et al., 2019) and  (Prillo and Eisenschlos, 2020). Those too are $O(N^{2})$ complexity solutions capable of ranking listings, but their focus is to replace traditional sorting algorithms like Quicksort with a neural network. In contrast, the goal of all-pairwise LTR is to improve ranking by learning interactions between listings, while still using Quicksort or Mergesort to perform the final sorting.
A frequent stumbling block is the apparent resemblance between all-pairwise LTR and its distant cousin, listwise LTR (Cao et al., 2007), a univariate model that optimizes the softmax cross-entropy loss. Since listwise LTR is univariate, it cannot account for interactions between listings. Similarly, methods that directly optimize NDCG  (Qin et al., 2010)  (Jagerman et al., 2022) also rely on univariate models, leaving them unable to model listing-to-listing interactions.
@@ 6. Optimizing Multiple Objectives
A popular solution to the problem is multi-task learning (Ren et al., 2023) (Tan et al., 2023), where a model predicts multiple objectives through different heads and combines them using weights. This approach makes the model more complex, since it has to handle multiple labels and balance dense and sparse ones. Particularly in our case, true-pairwise and all-pairwise LTR fit only one of the objectives, the booking probability, as searchers naturally compare various listings in search results. These methods aren’t suitable for predicting trip ratings, requiring a separate model to handle that task.
@@
We discard the multi-task and multi-model approaches to solve the multiple objectives puzzle. Instead, we unify the objectives by adjusting the training loss. For each training example consisting of a pair of listings $l_{\text{booked}}$ and $l_{\text{notBooked}}$, we compute a weight:
(10) $\omega_{\text{total}}=1+\omega_{\text{tripQuality}}(l_{\text{booked}})$
where a booking gets a base weight of $1$, and trip quality adds to it depending on the reviews of $l_{booked}$. The loss is calculated as:
(11) $\text{loss}=-\log(\theta(f(l_{\text{booked}})-f(l_{\text{notBooked}})))*\omega_{\text{total}}$